OpenStudy (anonymous):

James has two porcelain spheres. The radii of the spheres are in a ratio of 2 : 5. What is the ratio between the volumes of the two spheres?

OpenStudy (anonymous):

lets say the radius of the smaller sphere is 2 and the radius of the larger sphere is 5. the forumla for the volume of sa sphere is \[4/3(\pi*r^2)\] to find a ratio you can set up a fracgtion \[(4/3*\pi*2^3)/(4/3*\pi*5^3)\] \[2^3/5^3\] \[8/125\]

OpenStudy (anonymous):

Good explanation! thanks!